Find all real values of x for which
f(x) = 0.

f(x) = x3 − x

Answer by Alan3354(69443)

Find all real values of x for which
f(x) = 0.

f(x) = x3 − x
-------------x
If you mean x^3 - x = 0
---
x(x^2 - 1) = 0
x = 0, -1, +1